OpenStudy (anonymous):

Compare and contrast, in complete sentences, the difference between the products of a2a4 and (a2)4. Provide the expanded form and product of each. Help, please and thank you!!

OpenStudy (phoenixfire):

Well the first one is: \[a^2a^4=a^{2+4}=a^6\] And the second is: \[(a^2)^4=a^{(2)(4)}=a^8\] In one you add the exponents together while in the other you multiply them. As you can clearly see a^6 is not the equal to a^8.

OpenStudy (anonymous):

Thank you so much!!